# RAG vs. GraphRAG: A Systematic Evaluation and Key Insights

### v1

@@ 4.4 Improving QA Performance
We conduct experiments to verify the effectiveness of the two proposed strategies. Specifically, we evaluate overall performance across all selected datasets. For the MultiHop-RAG and NovelQA datasets, we use the overall accuracy, while for the NQ and HotPotQA datasets, we use the F1 score as the evaluation metric. The results are shown in Figure 3. From these results, we observe that both strategies generally enhance overall performance. For example, on the MultiHop-RAG dataset with Llama 3.1-70B, Selection and Integration improve the best method by 1.1% and 6.4%, respectively. When comparing the Selection and Integration strategies, the Integration strategy usually achieves higher performance than the Selection strategy. However, the Selection strategy processes each query using either RAG or GraphRAG, making it more efficient. In contrast, the Integration strategy yields better performance but requires each query to be processed by both RAG and GraphRAG, increasing computational cost.
@@ 5.1 Datasets and Evaluation Metrics
We adopt two widely used single-document query-based summarization datasets, SQuALITY Wang et al. (2022) and QMSum Zhong et al. (2021), and two multi-document query-based summarization datasets, ODSum-story and ODSum-meeting Zhou et al. (2023), for our evaluation. Unlike the LLM-generated global queries used in the unreleased datasets of Edge et al. (2024), most queries in the selected datasets focus on specific roles or events. Since these datasets contain one or more ground truth summaries for each query, we use ROUGE-2 Lin (2004) and BERTScore Zhang et al. (2019) as evaluation metrics to measure lexical and semantic similarity between the predicted and ground truth summaries.
@@ 5.2 Summarization Experimental Results
    RAG generally performs well on query-based summarization tasks. This is particularly true on multi-document summarization datasets, where they are often the best method.
@@
    KG-based GraphRAG benefit from combining triplets with their corresponding text. This improves performance by incorporating more details, making predictions closer to the ground truth summaries.
@@
    The Integration strategy is often comparable to RAG only performance. This strategy integrates retrieved content from both RAG and Community-GraphRAG (Local), resulting in performance similar to RAG alone.
@@ 5.3 Position Bias in Existing Evaluation
We further conduct an evaluation following Edge et al. (2024), using the LLM-as-a-Judge method to compare RAG and Community-based GraphRAG from two perspectives: Comprehensiveness and Diversity. Comprehensiveness focuses on detail, addressing the question: "How much detail does the answer provide to cover all aspects and details of the question?" Meanwhile, Diversity emphasizes global information, evaluating "Does the answer provide a broad and globally inclusive perspective?". The prompt and details are shown in Appendix A.5. Specifically, we input the summaries generated by RAG and GraphRAG into the prompt and ask the LLM to select the better one for each metric, following Edge et al. (2024). Additionally, to better account for the order in which the summaries are presented, we consider two scenarios. Order 1 (O1): We place the RAG summary appears before the GraphRAG summary and Order 2 (O2): GraphRAG appears before RAG. We compare the proportion of selected best samples from RAG and GraphRAG, where a higher proportion indicates better performance as predicted by the LLM.
The results of RAG vs. GraphRAG (Local) and RAG vs. GraphRAG (Global) on the QMSum and ODSum-story datasets are presented in Figure 4. More result can be found in Appendix A.6. We can make the following observations: (1) Position bias Shi et al. (2024); Wang et al. (2024) is evident in the LLM-as-a-Judge evaluations for summarization task, as changing the order of the two methods significantly affects the predictions. This effect is particularly strong in the comparison between RAG and GraphRAG (Local), where the LLMs make completely opposite decisions depending on the order, as shown in Figures 4(a) and 4(c). However, (2) Comparison between RAG and GraphRAG (Global): While the proportions vary, RAG consistently outperforms GraphRAG (Global) in Comprehensiveness but underperforms in Diversity as shown in Figures 4(b) and 4(d). This result suggests that Community-based GraphRAG with Global Search focuses more on the global aspects of whole corpus, whereas RAG captures more detailed information.
@@ 6 Conclusion
In this paper, we systematically evaluate and compare RAG and GraphRAG on general text-based tasks. Our analysis reveals the distinct strengths of RAG and GraphRAG in QA and query-based summarization, as well as evaluation challenges in summarization tasks, providing valuable insights for future research. Building on these findings, we propose two strategies to enhance QA performance. Future work can explore improving GraphRAG through better graph construction or developing novel approaches to combine RAG and GraphRAG methods for both effectiveness and efficiency.
@@
In this paper, we evaluate and compare RAG and GraphRAG on Question Answering and Query-based Summarization tasks. Future work can extend this study to additional tasks to further assess the strengths and applicability of GraphRAG. Additionally, the graph construction in all GraphRAG methods explored in this work relies on LLM-based construction, where LLMs extract entities and relations. However, other graph construction models designed for text processing exist and can be investigated in future studies. Finally, we primarily evaluate generation performance using Llama 3.1-8B-Instruct and Llama 3.1-70B-Instruct. Future research can explore other generation models for a more comprehensive comparison.
@@ A.1.1 Question Answering
  • NovelQA Tang and Yang (2024): NovelQA is a benchmark designed to evaluate the long-text understanding and retrieval ability of LLMs using manually curated questions about English novels exceeding 50,000 words. The dataset includes queries that focus on minor details or require cross-chapter reasoning, making them inherently challenging for LLMs. It covers various query types such as details, multi-hop, single-hop, character, meaning, plot, relation, setting, span, and times. Key challenges highlighted by NovelQA include grasping abstract meanings (meaning questions), understanding nuanced relationships (relation questions), and tracking temporal sequences and spatial extents (span and time questions), emphasizing the difficulty of maintaining and applying contextual information across long narratives. We use it for single-document QA task.
@@ A.2 More results on NovelQA dataset
A.3 RAG vs. GraphRAG Selection
@@
A.4 Query-based Summarization Results with Llama3.1-70B model
@@ A.5 The LLM-as-a-Judge Prompt
A.6 The LLM-as-a-Judge Results on more datasets

### v2

@@ 4.4 Improving QA Performance
We conduct experiments to verify the effectiveness of the two proposed strategies. Specifically, we evaluate overall performance across all selected datasets. For the MultiHop-RAG and NovelQA datasets, we use the overall accuracy, while for the NQ and HotPotQA datasets, we use the F1 score as the evaluation metric. The results are shown in Figure 3 and Appendix A.7. From these results, we observe that both strategies generally enhance overall performance. For example, on the MultiHop-RAG dataset with Llama 3.1-70B, Selection and Integration improve the best method by 1.1% and 6.4%, respectively. When comparing the Selection and Integration strategies, the Integration strategy usually achieves higher performance than the Selection strategy. However, Selection strategy processes each query using either RAG or GraphRAG, making it more efficient. In contrast, Integration strategy yields better performance but requires each query to be processed by both RAG and GraphRAG, increasing computational cost.
@@ 5.1 Datasets and Evaluation Metrics
We adopt two widely used single-document query-based summarization datasets, SQuALITY Wang et al. (2022) and QMSum Zhong et al. (2021), and two multi-document query-based summarization datasets, ODSum-story and ODSum-meeting Zhou et al. (2023), for our evaluation. Unlike the LLM-generated global queries used in the unreleased datasets of Edge et al. (2024), most queries in the selected datasets focus on specific roles or events. Since these datasets contain one or more human-written ground truth summaries for each query, we use ROUGE-2 Lin (2004) and BERTScore Zhang et al. (2019) as evaluation metrics to measure lexical and semantic similarity between the predicted and ground truth summaries.
@@ 5.2 Summarization Experimental Results
    RAG and HippoRAG2 generally performs well on query-based summarization tasks, primarily because they retrieve original text chunks that are more closely aligned with ground truth.
@@
    KG-based GraphRAG benefit from combining triplets with their corresponding text. This improves performance by incorporating more details, making predictions closer to the human-written ground truth summaries.
@@
    The Integration strategy often performs comparably to RAG alone, as explained in Appendix A.7.
@@ 5.3 Position Bias in Existing Evaluation
We further conduct an evaluation following Edge et al. (2024), using the LLM-as-a-Judge method from two perspectives: Comprehensiveness and Diversity. Comprehensiveness measures how well the summary covers details of the query, while Diversity evaluates whether the answer provides a broad and globally inclusive perspective. Full prompt details are in Appendix A.9. Specifically, we input summaries from RAG and GraphRAG into the prompt and ask the LLM to choose the better one for each metric. To account for position bias, we evaluate two orderings: Order 1 (O1) places the RAG summary first, and Order 2 (O2) places GraphRAG first. We report the proportion of times each method is preferred.
The results of RAG vs. GraphRAG (Local) and RAG vs. GraphRAG (Global) on the QMSum and ODSum-story datasets are presented in Figure 4. More result can be found in Appendix A.10. We can make the following observations: (1) Position bias Shi et al. (2024); Wang et al. (2024b) is evident in the LLM-as-a-Judge evaluations for summarization task, as changing the order of the two methods significantly affects the predictions. This effect is particularly strong in the comparison between RAG and GraphRAG (Local), where the LLMs make completely opposite decisions depending on the order, as shown in Figures 4(a) and 4(c). However, (2) Comparison between RAG and GraphRAG (Global): While the proportions vary, RAG consistently outperforms GraphRAG (Global) in Comprehensiveness but underperforms in Diversity as shown in Figures 4(b) and 4(d). This result suggests that Community-based GraphRAG with Global Search focuses more on the global aspects of whole corpus, whereas RAG captures more detailed information.
In addition to performance comparisons, we also include a time analysis of indexing, retrieval, and generation, as well as token and storage analyses for both methods in Appendix A.12, providing further insights into their practical trade-offs.
@@ 6 Conclusion
This work provides the first unified evaluation of RAG and GraphRAG across both reasoning- and summarization-oriented tasks, revealing when and why explicit graph structures help large language models. Our analyses demonstrate clear task-dependent strengths, RAG excels at single-hop, detail-oriented retrieval, while GraphRAG shines in multi-hop reasoning and global summarization.
Building on these insights, we proposed two hybrid strategies, Selection and Integration, that dynamically combine RAG and GraphRAG, improving QA accuracy by up to +6.4 points on MultiHop-RAG. We further uncovered practical limitations, graph incompleteness and evaluation bias, that point to the next generation of graph-augmented LLMs: systems that can automatically construct, refine, and reason over graphs with efficiency and reliability.
@@
In this paper, we evaluate and compare RAG and GraphRAG on Question Answering and Query-based Summarization tasks. Future work can extend this study to additional tasks to further assess the strengths and applicability of GraphRAG. For example, tasks such as planning Wu et al. (2024) and mathematical reasoning Feng et al. (2021) have been shown to benefit from graph representations. However, the distinct advantages and limitations of RAG and GraphRAG in these settings remain to be systematically explored. Additionally, the graph construction in all GraphRAG methods explored in this work relies on LLM-based construction, where LLMs extract entities and relations. However, other graph construction models designed for text processing exist and can be investigated in future studies. Finally, we only use the basic retriever for RAG and GraphRAG. There are also other retrievers such as GNN-based retrievers and LLM-based retrievers for GraphRAG. We do not use GNN-based or LLM-based retrieval methods due to limitations in our setting. GNN-based retrievers require node-level supervision, which is unavailable in our dynamically constructed and often incomplete graphs. LLM-based retrievers typically rely on fixed relation types, whereas our graphs contain diverse and open-ended relations. To ensure a fair and consistent comparison between RAG and GraphRAG, we adopt a simple retrieval approach. Nonetheless, exploring how advanced retrieval strategies can be adapted to dynamically constructed KGs in GraphRAG is an interesting direction for future work.
@@ A.1.1 Question Answering
  • NovelQA Wang et al. (2024a): NovelQA is a benchmark designed to evaluate the long-text understanding and retrieval ability of LLMs using manually curated questions about English novels exceeding 50,000 words. The dataset includes queries that focus on minor details or require cross-chapter reasoning, making them inherently challenging for LLMs. It covers various query types such as details, multi-hop, single-hop, character, meaning, plot, relation, setting, span, and times. Key challenges highlighted by NovelQA include grasping abstract meanings (meaning questions), understanding nuanced relationships (relation questions), and tracking temporal sequences and spatial extents (span and time questions), emphasizing the difficulty of maintaining and applying contextual information across long narratives. We use it for single-document QA task.
@@ A.2 More results on NovelQA dataset
A.3 Retrieval accuracy of different methods
In this subsection, we compare the retrieval accuracy of various methods. Since there is no ground-truth label for the retrieval step, we evaluate effectiveness by reporting retrieval accuracy, defined as the proportion of cases where the ground-truth answer appears in the retrieved content. We conduct this evaluation on the Hotpot and NQ datasets.
Table 12: Retrieval accuracy of different methods on Hotpot and NQ datasets
Method Hotpot NQ
RAG 0.886 0.867
KG-GraphRAG(Triplets only) 0.392 0.3218
KG-GraphRAG(Triplets+Text) 0.698 0.615
Community-GraphRAG (Local) 0.6753 0.422
Community-GraphRAG (Global) 0.886 0.833
As shown in the Table 12, KG-GraphRAG (Triplets only) achieves relatively low retrieval accuracy, particularly on NQ. This is primarily due to the incompleteness of the constructed knowledge graphs—only 65.8% of answer entities exist in the HotpotQA KG, and 65.5% in the NQ KG. In contrast, Community-GraphRAG, which leverages community-level summarization, demonstrates significantly better retrieval performance.
These findings highlight several potential directions for improvement:
  1. 1.
    Enhancing KG construction to increase entity and relation coverage.
  2. 2.
    Combining structured graph information with raw text to improve retrieval robustness and completeness.
A.4 Case studies for the question answering task
In this section, we present examples where RAG fails but GraphRAG succeeds. In Case 1 (Figure 5), RAG fails because it does not retrieve all the relevant chunks required for answering the multi-hop reasoning question. KG-RAG also fails due to missing information during knowledge graph construction. However, Community-GraphRAG is able to answer correctly by leveraging community-level summarizations that capture the necessary context. In Case 2 (Figure 6), both KG-GraphRAG and Community-GraphRAG succeed because they capture the reasoning chain—either through explicit graph paths or through summarization within the same community. In contrast, RAG fails due to insufficient retrieved information.
Case 1 Question: Alexander Kerensky was defeated and destroyed by the Bolsheviks in the course of a civil war that ended when ? RAG: Lesser battles of the war continued on the periphery for two more years, and minor skirmishes with the remnants of the White forces in the Far East continued well into 1923. The war ended in 1923 Answer: 1923 KG-GraphRAG: Answer: “Insufficient Information.” Community-GraphRAG: Alexander Kerensky: Alexander Kerensky was defeated and destroyed by the Bolsheviks in the course of the Russian Civil War. Russian Civil War: The Russian Civil War was a significant multi-party conflict that took place in the former Russian Empire from November 1917 to October 1922, following the October Revolution. Answer: October 1922 Gold Answer: October 1922
Figure 5: Case 1 from Hotpot dataset.
Case 2 Question: Which British first-generation jet-powered medium bomber was used in the South West Pacific theatre of World War II? RAG: Answer: “Insufficient Information.” KG-GraphRAG: No. 12 Squadron → COMBAT_IN → South West Pacific theatre of WWII; No. 2 Squadron → EQUIPPED_WITH → English Electric Canberra jets; Answer: English Electric Canberra jets Community-GraphRAG: English Electric Canberra: The English Electric Canberra is a British first-generation jet-powered medium bomber that was manufactured during the 1950s. …It saw action as a bomber unit in the South West Pacific theatre of World War II and, equipped with English Electric Canberra jets… Answer: English Electric Canberra Gold Answer: English Electric Canberra
Figure 6: Case 2 from Hotpot dataset.
A.5 Multi-step Retrieval
Multi-step, or iterative, retrieval Trivedi et al. (2022); Santhanam et al. (2021) is a widely adopted technique for enabling RAG to handle multi-step reasoning tasks. Specifically, at each iteration, new queries are generated based on the retrieval results from the previous step. The system then performs another round of retrieval using the updated queries, repeating the process until the problem is resolved or a predefined maximum number of iterations is reached. To further compare the performance of RAG and GraphRAG (Community-GraphRAG (Local)), we apply the multi-step retrieval method, specifically IRCoT Trivedi et al. (2022), to both approaches. Specifically, we select NQ, Hotpot, and MultiHop-RAG datasets with the Llama 3.1-70B model.
The results are presented in Table 13 and Table 14. We observe consistent trends with single-step retrieval: while multi-step retrieval improves the performance of both RAG and GraphRAG, their respective strengths remain. RAG continues to perform better on single-hop and detail-oriented questions, whereas GraphRAG outperforms on multi-hop and reasoning-intensive queries.
Table 13: Performance comparison of RAG and GraphRAG on NQ and hotpot dataset under the multi-step retrieval setting.
Dataset NQ Hotpot
Method P R F1 P R F1
RAG 74.55 67.82 68.18 66.34 63.99 63.88
GraphRAG 71.27 65.46 65.44 67.20 64.89 64.60
IRCOT + RAG 77.30 70.86 71.11 66.61 63.85 63.81
IRCOT + GraphRAG 73.82 67.57 67.51 88.60 69.01 66.31
Table 14: Performance comparison of RAG and GraphRAG on MultiHop-RAG dataset under the multi-step retrieval setting.
Method Inference Comparision Null Temporal Overall
RAG 94.85 56.31 91.36 25.73 65.77
GraphRAG 92.03 60.16 88.7 49.06 71.17
IRCoT+RAG 96.2 64.95 80.07 57.63 75.04
IRCoT+GraphRAG 94.98 65.89 69.44 60.38 74.33
A.6 RAG vs. GraphRAG Selection
@@ A.5 Multi-step Retrieval
A.7 RAG and GraphRAG Integration
In this section, we explore the effect of integrating RAG and GraphRAG for the question answering task. Specifically, we concatenate the retrieved results from both RAG and GraphRAG before passing them to the LLM. The results are presented in Table 15, Table 16, Table 17, Table 18, and Table 19, respectively. For most cases, the integration of RAG and GraphRAG improves performance. However, we observe a performance drop when integrating with Llama 3.1–8B on the MultiHop-RAG dataset. This degradation is primarily attributed to a significant decline on Null queries—those requiring the model to respond with “Insufficient Information.” By concatenating the retrieved results from both RAG and GraphRAG, the input length increases considerably, making the 8B model more susceptible to hallucination and the generation of incorrect answers. This vulnerability is more pronounced in the 8B model due to its limited capacity, whereas the 70B model demonstrates greater robustness to longer contexts and handles ambiguous information more conservatively. In contrast, for other query types such as Comparison and Temporal, the integration strategy yields notable gains on both model sizes.
Table 15: Performance comparison of RAG, GraphRAG, and their integration on NQ and Hotpot datasets
NQ Hotpot
Datasets Llama 3.1-8B Llama 3.1-70B Llama 3.1-8B Llama 3.1-70B
Method P R F1 P R F1 P R F1 P R F1
RAG 71.70 63.93 64.78 74.55 67.82 68.18 62.32 60.47 60.04 66.34 63.99 63.88
GraphRAG 69.48 62.54 63.01 71.27 65.46 65.44 64.14 62.08 61.66 67.20 64.89 64.60
Integration 72.81 65.91 66.28 75.67 69.75 69.75 67.21 65.09 64.76 69.22 66.70 66.50
Table 16: The performance of Llama 3.1-8B on MultiHop-RAG dataset
8B Inference Comparison Null Temporal Overall
RAG 92.16 57.59 96.01 30.7 67.02
GraphRAG 86.89 60.63 80.07 50.6 69.01
Integration 89.71 64.14 50.17 53.34 68.19
Table 17: The performance of Llama 3.1-70B on MultiHop-RAG dataset
70B Inference Comparison Null Temporal Overall
RAG 94.85 56.31 91.36 25.73 65.77
GraphRAG 92.03 60.16 88.7 49.06 71.17
Integration 96.45 73.48 59.47 66.72 77.62
Table 18: Performance of integrating RAG and GraphRAG with Llama 3.1–8B on the NovelQA dataset.
Integration character meaning plot relat settg span times avg
mh 70.83 58.82 63.89 73.03 84.62 60.00 36.90 49.17
sh 62.50 64.29 74.90 62.50 79.38 - - 70.85
dtl 60.00 43.82 83.93 21.43 72.97 - - 54.20
avg 63.33 50.19 75.23 60.80 78.23 60.00 36.90 58.36
For the query-based summarization task, we observed that the Integration strategy generally performs comparably to RAG, but not significantly better. This is because the evaluation is based on human-written ground-truth summaries, which tend to focus on detailed and faithful representations of the original text. RAG directly retrieves text segments that often match these detailed references more closely, as shown in Figure 4 of our paper. In contrast, GraphRAG primarily retrieves structured information (e.g., entities and relations), which omit finer details needed to align with ground-truth summaries. As a result, while Integration combines complementary views, the added structured content from GraphRAG does not consistently enhance alignment with detailed ground-truth summaries, leading to comparable or slightly lower scores.
Table 19: Performance of integrating RAG and GraphRAG with Llama 3.1–70B on the NovelQA dataset.
Integration character meaning plot relat settg span times avg
mh 77.08 70.59 83.33 77.53 92.31 44.00 37.97 51.99
sh 74.34 74.29 82.43 75.00 87.63 - - 80.04
dtl 67.14 53.37 92.86 75.00 89.19 - - 67.21
avg 72.96 60.00 84.29 76.80 88.44 44.00 37.97 65.97
A.8 Query-based Summarization Results with Llama3.1-70B model
@@ A.9 The LLM-as-a-Judge Prompt
A.10 The LLM-as-a-Judge Results on more datasets
@@
A.11 Graph Construction with different LLMs
In the main paper, we use GPT-4o-mini to extract entities and relationships for graph construction due to cost considerations. To investigate whether stronger LLMs yield better performance, we also use GPT-4o for graph extraction. Specifically, we evaluate this on the MultiHop-RAG and ODSum-story datasets, representing question answering and summarization tasks, respectively. We focused on Community-GraphRAG (Local) as a representative method (GraphRAG) and evaluated it with both LLaMA3.1-8B and LLaMA3.1-70B for generation.
The results are shown in Table 22, Table 23, Table 24 and Table 25, respectively. The results show that using a stronger LLM (GPT-4o) for graph extraction generally improves the performance of GraphRAG on both question answering and summarization tasks. However, the overall conclusion regarding the relative performance of RAG and GraphRAG remains consistent across different graph construction backbones.
Table 22: Performance of different graph construction methods with Llama 3.1–8B on the MultiHop-RAG dataset.
Inference Comparison NULL Temporal Overall
RAG 92.16 57.59 96.01 30.7 67.02
GPT-4o-mini 86.89 60.63 80.07 50.6 69.01
GPT-4o 88.11 62.62 70.43 49.74 68.74
Table 23: Performance of different graph construction methods with Llama 3.1–70B on the MultiHop-RAG dataset.
70B Inference Comparison NULL Temporal Overall
RAG 94.85 56.31 91.36 25.73 65.77
GPT-4o-mini 92.03 60.16 88.70 49.06 71.17
GPT-4o 93.63 66.59 81.06 58.49 75.08
Table 24: Performance of different graph construction methods with Llama 3.1–8B on the ODSum-story dataset.
ROUGE-2 BERTScore
P R F1 P R F1
RAG 15.39 8.44 9.81 83.87 85.74 84.57
GPT-4o-mini 13.84 7.19 8.49 83.19 85.07 83.90
GPT-4o 13.99 7.45 8.64 83.24 85.1 83.94
Table 25: Performance of different graph construction methods with Llama 3.1–8B on the ODSum-story dataset.
ROUGE-2 BERTScore
P R F1 P R F1
RAG 11.85 14.24 11.09 85.96 85.76 85.67
GPT-4o-mini 12.54 10.31 9.61 84.51 85.33 84.71
GPT-4o 12.08 10.84 9.72 84.66 85.28 84.77
A.12 Computation and Storage Analysis
In this section, we explore the computational and storage trade-offs of RAG, KG-GraphRAG, and Community-GraphRAG. We report construction time, retrieval time, and storage size on two representative datasets: MultiHop-RAG (for question answering) and ODSum-story (for summarization). The results are presented in the Table 26 and Table 27, respectively.
Table 26: The time and storage analysis on MultiHop-RAG dataset.
Method Construction Time Retrieval time Storage
RAG 135 1724 127MB
KG-GraphRAG 7702 14434 117MB
Community-GraphRAG 5560 1249 165MB
Table 27: The time and storage analysis on ODSum-story dataset
Method Construction Time Retrieval time Storage
RAG 74 350 71MB
KG-GraphRAG 6496 3527 44MB
Community-GraphRAG 2828 282 104MB
From the results, we have the following observations:
  • Construction time: KG-GraphRAG incurs the highest construction time due to the use of LLMs for triplet extraction. However, this process is performed offline.
  • Storage: GraphRAG variants generally consume less storage than RAG, with KG-GraphRAG being the most compact due to structured representations.
  • Retrieval time: KG-GraphRAG shows the highest latency, caused by LLM-based entity expansion during graph traversal. In contrast, Community-GraphRAG achieves the fastest retrieval through direct entity matching, even outperforming RAG.
We also assessed the average retrieved token count, retrieval time, generation time, and performance of our hybrid strategies (Selection and Integration) on the MultiHop-RAG dataset with the Llama3.1-70B model. The results are summarized in Table 28. As shown, the Integration strategy yields the highest performance but introduces the most overhead in terms of tokens and latency due to combining both RAG and GraphRAG content. In contrast, the Selection strategy provides a more balanced trade-off, improving performance over both RAG and GraphRAG individually, while keeping token and time costs significantly lower than GraphRAG.
Table 28: Comparison of retrieved tokens, retrieval time, generation time, and performance using Llama 3.1–70B.
Average Retrieved Tokens Retrieval Time Generation Time Performance
RAG 3631 1724 3640 65.77
GraphRAG 9770 1249 6272 71.17
Selection 8040 1562 5530 72.3
Integration 13401 2973 9674 77.62
Besides runtime and storage, we also analyze the number of tokens retrieved by Community-GraphRAG and RAG. The results are shown in Table 29.
Table 29: The retrieved number of tokens.
RAG Community-GraphRAG
MultiHop-RAG 3631 9770
ODSum-Story 2279 10244
In our experimental setup, RAG retrieves the top-10 text chunks, while Community-GraphRAG (Local) retrieves the top-10 entities and their associated relations. As shown in Table 29, Community-GraphRAG results in significantly more input tokens due to the inclusion of entities, entity descriptions, relations, relation descriptions, and community summaries.
To ensure a fair comparison, we conducted an additional experiment in which we increased the number of retrieved text chunks for RAG to match the total number of input tokens retrieved by Community-GraphRAG. The results are shown in Table 30, Table 31, Table 32 and Table 33. While increasing RAG’s input size does lead to slight performance gains, our main conclusions remain unchanged: RAG performs better on inference-style queries and summarization tasks, where detailed information is directly retrievable. In contrast, GraphRAG performs better on complex queries such as Comparison and Temporal types in MultiHop-RAG, which require multi-hop reasoning and aggregation.
Table 30: Performance comparison of RAG, token-matched RAG, and GraphRAG using Llama 3.1–8B on MultiHop-RAG dataset.
Inference Comparison NULL Temporal Overall
RAG 92.16 57.59 96.01 30.7 67.02
RAG_Same Token 95.34 59.81 89.04 36.71 69.33
GraphRAG 86.89 60.63 80.07 50.6 69.01
Table 31: Performance comparison of RAG, token-matched RAG, and GraphRAG using Llama 3.1–70B on MultiHop-RAG dataset.
70B Inference Comparison NULL Temporal Overall
RAG 94.85 56.31 91.36 25.73 65.77
RAG_Same Token 95.96 59.58 88.7 43.74 71.01
GraphRAG 92.03 60.16 88.7 49.06 71.17
Table 32: Performance comparison of RAG, token-matched RAG, and GraphRAG using Llama 3.1–8B on ODSum-Story dataset.
8B ROUGE-2 BERTScore
P R F1 P R F1
RAG 15.39 8.44 9.81 83.87 85.74 84.57
RAG_Same Token 14.16 10.02 10.16 84.34 85.74 84.82
GraphRAG 13.84 7.19 8.49 83.19 85.07 83.9
Table 33: Performance comparison of RAG, token-matched RAG, and GraphRAG using Llama 3.1–70B on ODSum-Story dataset.
ROUGE-2 BERTScore
P R F1 P R F1
RAG 11.85 14.24 11.09 85.96 85.76 85.67
RAG_Same Token 12.82 14.07 11.34 85.86 86 85.73
GraphRAG 12.54 10.31 9.61 84.51 85.33 84.71